# A solution to the Cauchy dual subnormality problem for a cyclic analytic $2$-isometry with defect operator of rank two

### v1

@@ 2. Case (ii)
Case 1: $\cos\theta<0.6$
Note that if $\Im(\alpha_{1}\overline{\alpha_{2}})=0$ then $\alpha_{1}\overline{\alpha_{2}}\in\mathbb{R}$, giving $\frac{\alpha_{1}}{\alpha_{2}}=\frac{\overline{\alpha_{1}}}{\overline{\alpha_{2}}}.$ This implies that $\frac{\alpha_{1}}{\alpha_{2}}$ is a real number, which is contradiction. Thus, $\Im(\alpha_{1}\overline{\alpha_{2}})\neq 0.$
Also, by Lemma 3.5, we know that $\alpha_{2}=\xi\overline{\alpha_{1}}$. Thus, we have, $\alpha_{1}\overline{\alpha_{2}}\in\mathbb{C}\setminus[1,\infty)$ and $\alpha_{2}\overline{\alpha_{1}}\in\mathbb{C}\setminus[1,\infty)$. Hence, the conditions to apply Corollary 3.3 are satisfied.
@@
After canceling the terms, we get
$p_{1}(\xi\overline{\alpha_{2}})\overline{p_{1}(\alpha_{2})}+p_{2}(\xi\overline{\alpha_{2}})\overline{p_{2}(\alpha_{2})}=\xi\overline{\alpha_{2}}^{2}(a_{11}+\xi a_{22}\overline{\alpha_{2}}^{2}).$
Finally, since $a_{11},a_{22}\in\mathbb{R}$ and $\xi\overline{\alpha_{2}}^{2}=\alpha_{1}\overline{\alpha_{2}}\in\mathbb{C}\setminus[1,\infty)$, we conclude that
$a_{11}+\xi a_{22}\overline{\alpha_{2}}^{2}\neq 0$
which implies
$p_{1}(\alpha_{1})\overline{p_{1}(\alpha_{2})}+p_{2}(\alpha_{1})\overline{p_{2}(\alpha_{2})}\neq 0.$
Hence, by Corollary 3.3, the Cauchy dual of operator $M_{z}$ is not subnormal.
Case 2: $\cos\theta>0.6$
We have already seen in Lemma 3.5 that whenever $\cos\theta>0.6$ we have $\arg(\alpha_{1})=\arg(\alpha_{2})$. We fix the notations for $\alpha_{1}$ and $\alpha_{2}$ such that $|\alpha_{1}|<|\alpha_{2}|$, in which case, we have $\alpha_{2}=k\alpha_{1}$ for some $k>1$. We use Theorem 3.2 to prove that the Cauchy dual of $M_{z}$ on $D(\mu)$ is not subnormal, where $\mu$ is a finite, positive Borel measure on $\mathbb{T}$ supported at $\{1,\xi\}$. We prove that for $l=1$, the $2\times 2$ minor of the infinite matrix (in Theorem 3.2)
@@
$k_{rt}=\displaystyle\frac{1}{a_{r}\overline{a_{t}}}\left(\sum_{j=1}^{2}p_{j}(\alpha_{r})\overline{p_{j}(\alpha_{t})}\right)\left(1-\frac{1}{\alpha_{r}\overline{\alpha_{t}}}\right)$
For $l=1$, we denote the $2\times 2$ principal minor as $M=\begin{vmatrix}m_{11}&m_{12}\\ m_{21}&m_{22}\end{vmatrix}$.
@@
$m_{11}$ $=\displaystyle\frac{k_{11}}{|\alpha_{1}|^{4}}+\frac{k_{12}}{k^{2}|\alpha_{1}|^{4}}+\frac{k_{21}}{k^{2}|\alpha_{1}|^{4}}+\frac{k_{22}}{k^{4}|\alpha_{1}|^{4}}$
$m_{12}$ $=\displaystyle\frac{k_{11}}{\overline{\alpha_{1}}|\alpha_{1}|^{4}}+\frac{k_{12}}{k^{3}\overline{\alpha_{1}}|\alpha_{1}|^{4}}+\frac{k_{21}}{k^{2}\overline{\alpha_{1}}|\alpha_{1}|^{4}}+\frac{k_{22}}{k^{5}\overline{\alpha_{1}}|\alpha_{1}|^{4}}$
$m_{21}$ $=\displaystyle\frac{k_{11}}{\alpha_{1}|\alpha_{1}|^{4}}+\frac{k_{12}}{k^{2}\alpha_{1}|\alpha_{1}|^{4}}+\frac{k_{21}}{k^{3}\alpha_{1}|\alpha_{1}|^{4}}+\frac{k_{22}}{k^{5}\alpha_{1}|\alpha_{1}|^{4}}$
$m_{22}$ $=\displaystyle\frac{k_{11}}{|\alpha_{1}|^{6}}+\frac{k_{12}}{k^{3}|\alpha_{1}|^{6}}+\frac{k_{21}}{k^{3}|\alpha_{1}|^{6}}+\frac{k_{22}}{k^{6}|\alpha_{1}|^{6}}$
Thus, we get,
$M=\displaystyle\frac{1}{|\alpha_{1}|^{10}}\Bigg[$ $k_{11}^{2}+\displaystyle\frac{k_{11}k_{12}}{k^{3}}+\frac{k_{11}k_{21}}{k^{3}}+\frac{k_{11}k_{22}}{k^{6}}+\frac{k_{12}k_{11}}{k^{2}}+\frac{k_{12}^{2}}{k^{5}}+\frac{k_{12}k_{21}}{k^{5}}$
$+\frac{k_{12}k_{22}}{k^{8}}+\frac{k_{21}k_{11}}{k^{2}}+\frac{k_{21}k_{12}}{k^{5}}+\frac{k_{21}^{2}}{k^{5}}+\frac{k_{21}k_{22}}{k^{8}}+\frac{k_{22}k_{11}}{k^{4}}+\frac{k_{22}k_{12}}{k^{7}}$
$+\frac{k_{22}k_{21}}{k^{7}}+\frac{k_{22}^{2}}{k^{10}}-k_{11}^{2}-\frac{k_{11}k_{12}}{k^{2}}-\frac{k_{11}k_{21}}{k^{3}}-\frac{k_{11}k_{22}}{k^{5}}-\frac{k_{12}k_{11}}{k^{3}}$
$-\frac{k_{12}^{2}}{k^{5}}-\frac{k_{12}k_{21}}{k^{6}}-\frac{k_{12}k_{22}}{k^{8}}-\frac{k_{21}k_{11}}{k^{2}}-\frac{k_{21}k_{12}}{k^{4}}-\frac{k_{21}^{2}}{k^{5}}-\frac{k_{21}k_{22}}{k^{7}}$
$-\frac{k_{22}k_{11}}{k^{5}}-\frac{k_{22}k_{12}}{k^{7}}-\frac{k_{22}k_{21}}{k^{8}}-\frac{k_{22}^{2}}{k^{10}}\Bigg]$
@@
$-\frac{k_{12}k_{21}}{k^{4}}+\frac{k_{22}k_{11}}{k^{4}}-\frac{k_{22}k_{11}}{k^{5}}\Bigg]$
Therefore,
$M=\displaystyle\frac{\left(1-\frac{1}{k}\right)^{2}}{k^{4}|\alpha_{1}|^{10}}\left(k_{11}k_{22}-k_{12}k_{21}\right)$
We prove that $D=k_{11}k_{22}-k_{12}k_{21}<0$.
We note that $|a_{1}|^{2}=|\alpha_{1}|^{2}(k-1)^{2}$, $|a_{2}|^{2}=|\alpha_{1}|^{2}(k-1)^{2}$.
Also, $a_{1}\overline{a_{2}}=-|\alpha_{1}|^{2}(k-1)^{2}=a_{2}\overline{a_{1}}$.
Let us fix the notations as $x=p_{1}(\alpha_{1})$, $y=p_{2}(\alpha_{1})$ and $w=p_{1}(\alpha_{2})$. Note that $p_{2}(\alpha_{2})=k^{2}y$.
Thus we get,
$D=\displaystyle\frac{1}{|\alpha_{1}|^{4}(k-1)^{4}}$ $\Bigg[\left(|x|^{2}+|y|^{2}\right)\left(1-\frac{1}{|\alpha_{1}|^{2}}\right)\left(|w|^{2}+k^{4}|y|^{2}\right)\left(1-\frac{1}{k^{2}|\alpha_{1}|^{2}}\right)$
$-\left(x\overline{w}+|y|^{2}k^{2}\right)\left(1-\frac{1}{k|\alpha_{1}|^{2}}\right)\left(w\overline{x}+|y|^{2}k^{2}\right)\left(1-\frac{1}{k|\alpha_{1}|^{2}}\right)\Bigg]$
$D$ $=\displaystyle\frac{1}{|\alpha_{1}|^{4}(k-1)^{4}}\Bigg[\left(|x|^{2}|w|^{2}+k^{4}|x|^{2}|y|^{2}+|y|^{2}|w|^{2}+k^{4}|y|^{4}\right)\times$
$\hskip 113.81102pt\left(1-\frac{1}{|\alpha_{1}|^{2}}-\frac{1}{k^{2}|\alpha_{1}|^{2}}+\frac{1}{k^{2}|\alpha_{1}|^{4}}\right)$
$\hskip 42.67912pt-\left(|x|^{2}|w|^{2}+2k^{2}|y|^{2}\Re(x\overline{w})+|y|^{4}k^{4}\right)\left(1-\frac{2}{k|\alpha_{1}|^{2}}+\frac{1}{k^{2}|\alpha_{1}|^{4}}\right)\Bigg]$
$D=\displaystyle\frac{1}{|\alpha_{1}|^{4}(k-1)^{4}}\Bigg[$ $|x|^{2}|w|^{2}+k^{4}|x|^{2}|y|^{2}+|y|^{2}|w|^{2}+k^{4}|y|^{4}$
$-\frac{|x|^{2}|w|^{2}}{|\alpha_{1}|^{2}}-\frac{k^{4}|x|^{2}|y|^{2}}{|\alpha_{1}|^{2}}-\frac{|y|^{2}|w|^{2}}{|\alpha_{1}|^{2}}-\frac{k^{4}|y|^{4}}{|\alpha_{1}|^{2}}$
$-\frac{|x|^{2}|w|^{2}}{k^{2}|\alpha_{1}|^{2}}-\frac{k^{2}|x|^{2}|y|^{2}}{|\alpha_{1}|^{2}}-\frac{|y|^{2}|w|^{2}}{k^{2}|\alpha_{1}|^{2}}-\frac{k^{2}|y|^{4}}{|\alpha_{1}|^{2}}$
$+\frac{|x|^{2}|w|^{2}}{k^{2}|\alpha_{1}|^{4}}+\frac{k^{2}|x|^{2}|y|^{2}}{|\alpha_{1}|^{4}}+\frac{|y|^{2}|w|^{2}}{k^{2}|\alpha_{1}|^{4}}+\frac{k^{2}|y|^{4}}{|\alpha_{1}|^{4}}$
$-|x|^{2}|w|^{2}-2k^{2}|y|^{2}\Re(x\overline{w})-|y|^{4}k^{4}+\frac{2|x|^{2}|w|^{2}}{k|\alpha_{1}|^{2}}$
$+\frac{4k|y|^{2}\Re(x\overline{w})}{|\alpha_{1}|^{2}}+\frac{2k^{3}|y|^{4}}{|\alpha_{1}|^{2}}-\frac{|x|^{2}|w|^{2}}{k^{2}|\alpha_{1}|^{4}}$
$-\frac{2|y|^{2}\Re(x\overline{w})}{|\alpha_{1}|^{4}}-\frac{k^{2}|y|^{4}}{|\alpha_{1}|^{4}}\Bigg]$
Thus,
$D=\displaystyle\frac{1}{|\alpha_{1}|^{4}(k-1)^{4}}\Bigg[$ $\frac{|x|^{2}|w|^{2}}{|\alpha_{1}|^{2}}\left(-1-\frac{1}{k^{2}}+\frac{2}{k}\right)+\frac{k^{2}|x|^{2}|y|^{2}}{|\alpha_{1}|^{2}}\left(-k^{2}-1+\frac{1}{|\alpha_{1}|^{2}}\right)$
$+\frac{|y|^{2}|w|^{2}}{|\alpha_{1}|^{2}}\left(-1-\frac{1}{k^{2}}+\frac{1}{k^{2}|\alpha_{1}|^{2}}\right)+\frac{k^{2}|y|^{4}}{|\alpha_{1}|^{2}}\left(-k^{2}-1+2k\right)$
$+2|y|^{2}\Re(x\overline{w})\left(-k^{2}+\frac{2k}{|\alpha_{1}|^{2}}-\frac{1}{|\alpha_{1}|^{4}}\right)\Bigg]$
$D=\displaystyle\frac{1}{|\alpha_{1}|^{4}(k-1)^{4}}\Bigg[$ $-\frac{|x|^{2}|w|^{2}}{|\alpha_{1}|^{2}}\left(1-\frac{1}{k}\right)^{2}-\frac{k^{2}|x|^{2}|y|^{2}}{|\alpha_{1}|^{2}}\left(k^{2}+1-\frac{1}{|\alpha_{1}|^{2}}\right)$
@@
$-2|y|^{2}\Re(x\overline{w})\left(k-\frac{1}{|\alpha_{1}|^{2}}\right)^{2}\Bigg]$ (3.21)
We consider the expression:
$L=$ $\frac{k^{2}|x|^{2}|y|^{2}}{|\alpha_{1}|^{2}}\left(k^{2}+1-\frac{1}{|\alpha_{1}|^{2}}\right)+\frac{|y|^{2}|w|^{2}}{|\alpha_{1}|^{2}}\left(1+\frac{1}{k^{2}}-\frac{1}{k^{2}|\alpha_{1}|^{2}}\right)$
$+2|y|^{2}\Re(x\overline{w})\left(k-\frac{1}{|\alpha_{1}|^{2}}\right)^{2}$
$=$ $\frac{k^{2}|x|^{2}|p_{22}|^{2}|\alpha_{1}|^{4}}{|\alpha_{1}|^{2}}\left(\frac{|\alpha_{2}|^{2}+|\alpha_{1}|^{2}-1}{|\alpha_{1}|^{2}}\right)+\frac{|p_{22}|^{2}|\alpha_{1}|^{4}|w|^{2}}{|\alpha_{1}|^{2}}\left(\frac{|\alpha_{2}|^{2}+|\alpha_{1}|^{2}-1}{k^{2}|\alpha_{1}|^{2}}\right)$
$+2|p_{22}|^{2}|\alpha_{1}|^{4}\Re(x\overline{w})\frac{(|\alpha_{1}||\alpha_{2}|-1)^{2}}{|\alpha_{1}|^{4}}$ (3.22)
Let $\ell_{1}=|\alpha_{1}|$ and $\ell_{2}=|\alpha_{2}|$. Then we know that $\ell_{1},\ell_{2}>1$. Let $\alpha_{3},\alpha_{4}$ are the other two roots of $f(z)=z^{4}-3(1+\xi)z^{3}+(8+2\cos\theta)\xi z^{2}-3(\xi+\xi^{2})z+\xi^{2}$. Then we know that $\alpha_{2}=\frac{1}{\overline{\alpha_{1}}}$ and hence $|\alpha_{3}|,|\alpha_{4}|<1$.
We have,
$\sum_{i=1}^{4}\alpha_{i}=3(1+\xi)$
@@
But, $\cos\theta>0.6$ gives, $\sum_{i=1}^{4}|\alpha_{i}|\geq\displaystyle\frac{12}{\sqrt{5}}$.
Since, $|\alpha_{3}|,|\alpha_{4}|<1$, we get, $\ell_{1}+\ell_{2}>\displaystyle\frac{12}{\sqrt{5}}-2$.
Now, $\sqrt{5}<\frac{7}{3}$ and hence, $\frac{12}{\sqrt{5}}-2>3$.
Thus we get, $\ell_{1}+\ell_{2}>3$.
Also, we have, $(\ell_{1}-1)(\ell_{2}-1)>0$ which gives us
$\ell_{1}\ell_{2}>\ell_{1}+\ell_{2}-1>3-1=2.$
Thus, we get, $b=\ell_{1}\ell_{2}>2$ whenever $\cos\theta>0.6$.
Hence, $|\alpha_{1}||\alpha_{2}|-1>1$ which gives us
$\left(|\alpha_{1}||\alpha_{2}|-1\right)^{2}>|\alpha_{1}||\alpha_{2}|-1.$ (3.23)
Further, by AM-GM inequality, we get, $\ell_{1}^{2}+\ell_{2}^{2}\geq 2\ell_{1}\ell_{2}$.
Also, $2\ell_{1}\ell_{2}\geq\ell_{1}\ell_{2}$.
@@
$|\alpha_{1}|^{2}+|\alpha_{2}|^{2}-1\geq|\alpha_{1}||\alpha_{2}|-1.$ (3.24)
@@
$>\left(\frac{|\alpha_{1}\alpha_{2}|-1}{k^{2}}\right)|p_{22}|^{2}\left(|x|^{2}+|w|^{2}+2\Re(x\overline{w})\right)..\{\text{As }k>1,k^{2}\geq 1/k^{2},1\geq 1/k^{2}$
$=\left(\frac{|\alpha_{1}\alpha_{2}|-1}{k^{2}}\right)|p_{22}|^{2}\|x+w\|^{2}>0.$
Using this in the equation (3.21) we get $D<0$, which implies $M<0$. Thus, the Cauchy dual of $M_{z}$ is not subnormal.
Case 3: Suppose $\cos\theta=0.6$.
We now deal with the case when $\cos\theta=0.6$.
As noted in Part 3 of Lemma 3.5, when $\cos\theta=0.6$ we have
$\alpha_{1}=\alpha_{2}=2+i=\alpha(say).$
We apply Theorem 3.7 to obtain $B=(b_{1},b_{2})$ such that $D(\mu)=H(B)$ with equality of norms where $b_{j}=\frac{p_{j}}{q}$, for $j=1,2$. By equations (3.19) and (3.20), we have $p_{1}(z)=p_{11}z+p_{12}z^{2}$ and $p_{2}(z)=p_{22}z^{2}$, where,

### v2

@@ 2. Case (ii)
Case 1: Let $\cos\theta<0.6.$
Note that if $\Im(\alpha_{1}\overline{\alpha_{2}})=0$ then $\frac{\alpha_{1}}{\alpha_{2}}=\frac{\overline{\alpha_{1}}}{\overline{\alpha_{2}}}.$ This implies that $\frac{\alpha_{1}}{\alpha_{2}}$ is a real number, which is not true as $\arg(\alpha_{1})\neq\arg(\alpha_{2})$. Thus, $\Im(\alpha_{1}\overline{\alpha_{2}})\neq 0.$
Also, by Lemma 3.5, we know that $\alpha_{2}=\xi\overline{\alpha_{1}}$.
Thus, we have, $\alpha_{1}\overline{\alpha_{2}}\in\mathbb{C}\setminus[1,\infty)$ and $\alpha_{2}\overline{\alpha_{1}}\in\mathbb{C}\setminus[1,\infty)$.
@@
Simplifying, we get
$p_{1}(\xi\overline{\alpha_{2}})\overline{p_{1}(\alpha_{2})}+p_{2}(\xi\overline{\alpha_{2}})\overline{p_{2}(\alpha_{2})}=\xi\overline{\alpha_{2}}^{2}(a_{11}+(\overline{a_{12}}+a_{12}\xi)\overline{\alpha_{2}}+a_{22}\xi\overline{\alpha_{2}}^{2}).$
Let $z_{1}=a_{11}$, $z_{2}=(\overline{a_{12}}+a_{12}\xi)\overline{\alpha_{2}}$ and $z_{3}=a_{22}\xi\overline{\alpha_{2}}^{2}.$
As $z_{1}\in\mathbb{R}$, it is enough to prove that $\Im(z_{2}+z_{3})\neq 0.$
Without loss of generality, we assume $\alpha_{2}$ as that root of $f(z)$, lying outside the closed unit disc, whose argument is smaller. We denote $\arg(\alpha_{2})$ by $\theta_{2}$.
  1. Case (i)
    Let $\pi/2<\theta<\pi$. We have $z_{2}=-2b\cos(\theta)(1+\xi)\overline{\alpha_{2}}$ and hence, $\arg(z_{2})=\frac{\theta}{2}-\theta_{2}$ and $\arg(z_{3})=\theta-2\theta_{2}.$
    Recall that $\theta_{1}=\arg(\alpha_{1})$ and we have $\theta_{1}+\theta_{2}=\theta.$
    Further, we have $\arg(1+\xi)=\frac{\theta}{2}$ and $\alpha_{1}+\alpha_{2}=a(1+\xi).$ This implies that $\arg(\alpha_{1}+\alpha_{2})=\arg(1+\xi)=\frac{\theta}{2}.$
    Since $|\alpha_{1}|=|\alpha_{2}|$, $\alpha_{1}+\alpha_{2}$ is the angle bisector of the angle between $\alpha_{1},\alpha_{2}$ and thus we must have $0<\arg(\alpha_{1})-\arg(\alpha_{2})<\pi.$ That is, $0<\theta-2\theta_{2}<\pi.$ Thus, $0<\frac{\theta}{2}-\theta_{2}<\frac{\pi}{2}.$ Hence, both $z_{2}$ and $z_{3}$ lie in the upper half plane which imply that $\Im(z_{2}+z_{3})\neq 0.$
  2. Case (ii)
    If $t_{0}<\theta<\pi/2$ then we prove that $\Im(z_{2}+z_{3})\neq 0.$
    We have $z_{2}=-2b\cos(\theta)(1+\xi)\overline{\alpha_{2}}.$ Note that $\cos\theta>0$ and $\theta_{2}<\frac{\theta}{2}$ implies that $z_{2}$ lies in third quadrant. Hence, $\arg(z_{2})=\frac{\theta}{2}-\theta_{2}-\pi.$
    Now, $z_{3}=a_{22}\xi\overline{\alpha_{2}}^{2}$ and thus, $\arg(z_{3})=\theta-2\theta_{2}.$ Hence, $z_{3}$ lies in first quadrant.
    $\Im(z_{2})=|z_{2}|\sin(\frac{\theta}{2}-\theta_{2}-\pi)$ and $\Im(z_{3})=|z_{3}|\sin(\theta-2\theta_{2}).$
    Suppose $\Im(z_{2}+z_{3})=0$, then we get,
    $|z_{3}|\sin(\theta-2\theta_{2})=|z_{2}|\sin(\frac{\theta}{2}-\theta_{2}).$
    Thus,
    (3.21) $\left|\frac{z_{3}}{z_{2}}\right|=\frac{1}{2\cos(\frac{\theta}{2}-\theta_{2})}<\frac{1}{2\cos(\frac{\theta}{2})}<\frac{1}{\sqrt{2}}.$
    We have, $|z_{2}|=2b|1+\xi||\cos(\theta)||\alpha_{2}|$ and $|z_{3}|=a_{22}|\alpha_{2}|^{2}=a_{22}b.$
    Thus, $\left|\frac{z_{2}}{z_{3}}\right|=\frac{4\cos(\theta)\cos(\theta/2)|\alpha_{2}|}{a_{22}}.$
    We have, $\cos(\theta)<0.6,\cos(\theta/2)<\sqrt{\frac{1+\cos\theta}{2}}<0.8944$ and
    $|\alpha_{2}|=\sqrt{b}<\sqrt{5.46269}=2.3372.$
    This gives, $4\cos(\theta)\cos(\theta/2)|\alpha_{2}|<5.0169.$
    We have, $a_{22}=1-3b+2ab-\frac{b}{\Delta}=1+b(2a-3-\frac{b}{\Delta}).$ We know that $5<b<6$. Now $a$ is an increasing function of $\theta$ (see part (2) of Lemma 3.7). This gives $2.5<a<2.5357.$ Further, $0.7<\Delta$ (see part (3) of Lemma 3.7). Thus we get,
    $a_{22}=1+b(2a-3-\frac{b}{\Delta})>1+5(2-\frac{1}{0.7})=3.85714.$
    Thus, $\left|\frac{z_{3}}{z_{2}}\right|>0.7688>\frac{1}{\sqrt{2}}.$ This is a contradiction to (3.21).
    Hence, we must have $\Im(z_{2}+z_{3})\neq 0.$
In the case when $\pi<\theta<-t_{0}$, by the symmetry, it can be observed that $\overline{\alpha_{2}}$ is the root of $f(z)$ corresponding to $\theta.$ Now, by applying Case (i) and Case(ii) to $\theta$, we get, $\Im(\overline{z_{2}+z_{3}})\neq 0$ and thus $\Im(z_{2}+z_{3})\neq 0.$
Thus we get, $p_{1}(\alpha_{1})\overline{p_{1}(\alpha_{2})}+p_{2}(\alpha_{1})\overline{p_{2}(\alpha_{2})}\neq 0.$ Hence, by Corollary 3.3, the Cauchy dual of the operator $M_{z}$ is not subnormal.
Case 2: Let $\cos\theta>0.6.$
By Lemma 3.5 we know that if $\cos\theta>0.6$ then $\arg(\alpha_{1})=\arg(\alpha_{2})$ and $|\alpha_{1}|\neq|\alpha_{2}|$. Without loss, suppose $|\alpha_{1}|<|\alpha_{2}|$ and $\alpha_{2}=k\alpha_{1}$ for some $k>1$. Therefore, note that Corollary 3.3 is not applicable in this case. Here, we apply Theorem 3.2 to reach the desired conclusion. It is enough to prove that for $l=1$, the $2\times 2$ minor of the infinite matrix
$\sum_{r,t=1}^{2}\left(\frac{1}{a_{r}\overline{a_{t}}}\sum_{j=1}^{2}p_{j}(\alpha_{r})\overline{p_{j}(\alpha_{t})}\right)\left(1-\frac{1}{\alpha_{r}\overline{\alpha_{t}}}\right)^{l}\left(\left(\frac{1}{\alpha_{r}^{m+2}\overline{\alpha_{t}}^{n+2}}\right)\right)_{m,n\geq 0}$
@@
(3.22) $k_{rt}=\displaystyle\frac{1}{a_{r}\overline{a_{t}}}\left(\sum_{j=1}^{2}p_{j}(\alpha_{r})\overline{p_{j}(\alpha_{t})}\right)\left(1-\frac{1}{\alpha_{r}\overline{\alpha_{t}}}\right).$
For $l=1$, we denote the $2\times 2$ principal minor as $M=\begin{vmatrix}m_{00}&m_{01}\\ m_{10}&m_{11}\end{vmatrix}$.
@@
$m_{00}$ $=\displaystyle\frac{k_{11}}{|\alpha_{1}|^{4}}+\frac{k_{12}}{k^{2}|\alpha_{1}|^{4}}+\frac{k_{21}}{k^{2}|\alpha_{1}|^{4}}+\frac{k_{22}}{k^{4}|\alpha_{1}|^{4}},$
$m_{01}$ $=\displaystyle\frac{k_{11}}{\overline{\alpha_{1}}|\alpha_{1}|^{4}}+\frac{k_{12}}{k^{3}\overline{\alpha_{1}}|\alpha_{1}|^{4}}+\frac{k_{21}}{k^{2}\overline{\alpha_{1}}|\alpha_{1}|^{4}}+\frac{k_{22}}{k^{5}\overline{\alpha_{1}}|\alpha_{1}|^{4}},$
$m_{10}$ $=\displaystyle\frac{k_{11}}{\alpha_{1}|\alpha_{1}|^{4}}+\frac{k_{12}}{k^{2}\alpha_{1}|\alpha_{1}|^{4}}+\frac{k_{21}}{k^{3}\alpha_{1}|\alpha_{1}|^{4}}+\frac{k_{22}}{k^{5}\alpha_{1}|\alpha_{1}|^{4}},$
$m_{11}$ $=\displaystyle\frac{k_{11}}{|\alpha_{1}|^{6}}+\frac{k_{12}}{k^{3}|\alpha_{1}|^{6}}+\frac{k_{21}}{k^{3}|\alpha_{1}|^{6}}+\frac{k_{22}}{k^{6}|\alpha_{1}|^{6}}.$
@@
$-\frac{k_{12}k_{21}}{k^{4}}+\frac{k_{22}k_{11}}{k^{4}}-\frac{k_{22}k_{11}}{k^{5}}\Bigg].$
Simplifying we get, $M=\displaystyle\frac{\left(1-\frac{1}{k}\right)^{2}}{k^{4}|\alpha_{1}|^{10}}\left(k_{11}k_{22}-k_{12}k_{21}\right).$
We denote the expression $k_{11}k_{22}-k_{12}k_{21}$ by $G$.
We prove that $G<0$. Note that $|a_{1}|^{2}=|\alpha_{1}|^{2}(k-1)^{2}$, $|a_{2}|^{2}=|\alpha_{1}|^{2}(k-1)^{2}$. Also, $a_{1}\overline{a_{2}}=-|\alpha_{1}|^{2}(k-1)^{2}=a_{2}\overline{a_{1}}$.
We introduce the following notations in order to achieve simplicity in the computations.
Let $x=p_{1}(\alpha_{1})$, $y=p_{2}(\alpha_{1})$ and $w=p_{1}(\alpha_{2})$. Note that $p_{2}(\alpha_{2})=k^{2}y$.
Substituting the values of $k_{11},k_{12},k_{21},k_{22}$ from equation (3.22) we get,
$G=\displaystyle\frac{1}{|\alpha_{1}|^{4}(k-1)^{4}}$ $\Bigg[\left(|x|^{2}+|y|^{2}\right)\left(1-\frac{1}{|\alpha_{1}|^{2}}\right)\left(|w|^{2}+k^{4}|y|^{2}\right)\left(1-\frac{1}{k^{2}|\alpha_{1}|^{2}}\right)$
$-\left(x\overline{w}+|y|^{2}k^{2}\right)\left(1-\frac{1}{k|\alpha_{1}|^{2}}\right)\left(w\overline{x}+|y|^{2}k^{2}\right)\left(1-\frac{1}{k|\alpha_{1}|^{2}}\right)\Bigg].$
Simplifying we get,
$G=\displaystyle\frac{1}{|\alpha_{1}|^{4}(k-1)^{4}}\Bigg[$ $-\frac{|x|^{2}|w|^{2}}{|\alpha_{1}|^{2}}\left(1-\frac{1}{k}\right)^{2}-\frac{k^{2}|x|^{2}|y|^{2}}{|\alpha_{1}|^{2}}\left(k^{2}+1-\frac{1}{|\alpha_{1}|^{2}}\right)$
@@
(3.23) $-2|y|^{2}\Re(x\overline{w})\left(k-\frac{1}{|\alpha_{1}|^{2}}\right)^{2}\Bigg]$
$=\displaystyle\frac{1}{|\alpha_{1}|^{4}(k-1)^{4}}\Bigg[-\frac{|x|^{2}|w|^{2}}{|\alpha_{1}|^{2}}\left(1-\frac{1}{k}\right)^{2}-\frac{k^{2}|y|^{4}}{|\alpha_{1}|^{2}}(k-1)^{2}-L\Bigg],$
where
$L=\frac{k^{2}|x|^{2}|y|^{2}}{|\alpha_{1}|^{2}}\left(k^{2}+1-\frac{1}{|\alpha_{1}|^{2}}\right)+\frac{|y|^{2}|w|^{2}}{|\alpha_{1}|^{2}}\left(1+\frac{1}{k^{2}}-\frac{1}{k^{2}|\alpha_{1}|^{2}}\right)+2|y|^{2}\Re(x\overline{w})\left(k-\frac{1}{|\alpha_{1}|^{2}}\right)^{2}.$
We now prove that $L>0$. This will in turn prove that $G<0$.
Substituting $y$ in $L$ we have
$L=$ $\frac{k^{2}|x|^{2}|p_{22}|^{2}|\alpha_{1}|^{4}}{|\alpha_{1}|^{2}}\left(\frac{|\alpha_{2}|^{2}+|\alpha_{1}|^{2}-1}{|\alpha_{1}|^{2}}\right)+\frac{|p_{22}|^{2}|\alpha_{1}|^{4}|w|^{2}}{|\alpha_{1}|^{2}}\left(\frac{|\alpha_{2}|^{2}+|\alpha_{1}|^{2}-1}{k^{2}|\alpha_{1}|^{2}}\right)$
(3.24) $+2|p_{22}|^{2}|\alpha_{1}|^{4}\Re(x\overline{w})\frac{(|\alpha_{1}||\alpha_{2}|-1)^{2}}{|\alpha_{1}|^{4}}.$
Let $\alpha_{3}=\frac{1}{\overline{\alpha_{1}}},\alpha_{4}=\frac{1}{\overline{\alpha_{2}}}$. Since $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ are the roots of
$f(z)=z^{4}-3(1+\xi)z^{3}+(8+2\cos\theta)\xi z^{2}-3(\xi+\xi^{2})z+\xi^{2},$
we get, $\sum_{i=1}^{4}\alpha_{i}=3(1+\xi).$
@@
But, $\cos\theta>0.6$ gives, $\sum_{i=1}^{4}|\alpha_{i}|\geq\displaystyle\frac{12}{\sqrt{5}}$. Since, $|\alpha_{3}|<1,|\alpha_{4}|<1$, we get, $|\alpha_{1}|+|\alpha_{2}|>\displaystyle\frac{12}{\sqrt{5}}-2$. Now, $\sqrt{5}<\frac{7}{3}$ and hence, $\frac{12}{\sqrt{5}}-2>3$. Thus we get, $|\alpha_{1}|+|\alpha_{2}|>3$. Also, we have, $(|\alpha_{1}|-1)(|\alpha_{2}|-1)>0$ which gives us
$|\alpha_{1}||\alpha_{2}|>|\alpha_{1}|+|\alpha_{2}|-1>3-1=2.$
Thus, we get, $b=|\alpha_{1}||\alpha_{2}|>2$ whenever $\cos\theta>0.6$. Hence, $|\alpha_{1}||\alpha_{2}|-1>1$ which gives us
(3.25) $\left(|\alpha_{1}||\alpha_{2}|-1\right)^{2}>|\alpha_{1}||\alpha_{2}|-1.$
Further, we have, $|\alpha_{1}|^{2}+|\alpha_{2}|^{2}\geq 2|\alpha_{1}||\alpha_{2}|$ and $2|\alpha_{1}||\alpha_{2}|\geq|\alpha_{1}||\alpha_{2}|$.
@@
(3.26) $|\alpha_{1}|^{2}+|\alpha_{2}|^{2}-1\geq|\alpha_{1}||\alpha_{2}|-1.$
@@
$=\left(|\alpha_{1}||\alpha_{2}|-1\right)|p_{22}|^{2}\left\|kx+\frac{w}{k}\right\|^{2}>0.$
Hence, we get $G<0$, which implies $M<0$. Thus, the Cauchy dual of $M_{z}$ on $D(\mu)$ is not subnormal.
Case 3: Suppose $\cos\theta=0.6$. As noted in Part (3) of Lemma 3.5, if $\cos\theta=0.6$, we have $\alpha_{1}=\alpha_{2}=2+i=\alpha(say).$
We apply Theorem 3.8 to obtain $B=(b_{1},b_{2})$ such that $D(\mu)=H(B)$ with the equality of norms where $b_{j}=\frac{p_{j}}{q}$, for $j=1,2$. By equations (3.19) and (3.20), we have $p_{1}(z)=p_{11}z+p_{12}z^{2}$ and $p_{2}(z)=p_{22}z^{2}$, where,